OpenStudy (anonymous):

Find the volume of the rectangular prism. A)4 m3 B)8 m3 C)7 m3 D)28 m3

OpenStudy (kamille):

what is given?

OpenStudy (kamille):

V=a*b*c a=4 b=2 c=1

OpenStudy (anonymous):

?

OpenStudy (anonymous):

so it 4

OpenStudy (anonymous):

lol

OpenStudy (kamille):

... V=4*2*1=?

OpenStudy (kamille):

*means multiplying

OpenStudy (anonymous):

oh lol

OpenStudy (anonymous):

ik

OpenStudy (anonymous):

what is v

OpenStudy (anonymous):

nvm

OpenStudy (kamille):

v is volume:)

OpenStudy (anonymous):

ik i got 8

OpenStudy (kamille):

yes,it is right aswer